the fractional measurement of 3 12/16 inches would equate to the decimal measurement of what?


1
Expert's answer
2021-01-21T11:58:10-0500

The fractional measurement of 3 12/16 inches would equate to the following decimal measurement:


"3\frac{12}{16}=3\frac{3}{4}=3+\frac{3}{4}=3.75''."